# Efficient Manipulation-Enhanced Semantic Mapping With Uncertainty-Informed Action Selection

### v1

@@ Abstract
Service robots operating in cluttered human environments such as homes, offices, and schools cannot rely on predefined object arrangements and must continuously update their semantic and spatial estimates while dealing with possible frequent rearrangements. Efficient and accurate mapping under such conditions demands selecting informative viewpoints and targeted manipulations to reduce occlusions and uncertainty. In this work, we present a manipulation-enhanced semantic mapping framework for occlusion-heavy shelf scenes that integrates evidential metric-semantic mapping with reinforcement-learning-based next-best view planning and targeted action selection. Our method thereby exploits uncertainty estimates from the Dirichlet and Beta distributions in the semantic and occupancy prediction networks to guide both active sensor placement and object manipulation, focusing on areas of limited knowledge and selecting actions with high expected information gain. For object manipulation, we introduce an uncertainty-informed push strategy that targets occlusion-critical objects and generates minimally invasive actions to reveal hidden regions. The experimental evaluation shows that our framework highly reduces object displacement and drops while achieving a 95% reduction in planning time compared to the state-of-the-art, thereby realizing real-world applicability.
@@ I Introduction
The successful deployment of general-purpose service robots in human-centric environments, such as homes and offices, depends on their ability to perceive and manipulate diverse objects within cluttered and constrained spaces. In such environments, robots must not only go beyond traditional passive perception approaches, such as static mapping, but also extend conventional active perception methods [1] by integrating physical interactions with the environment [2]. This allows them to actively remove occlusions by physically uncovering hidden regions [3, 4], which is especially critical for tasks such as exploring unknown environments [5], constructing detailed object models [6], and locating objects obscured from direct view [7]. However, such manipulation actions come at a cost, as uninformed physical interactions often not only increase the uncertainty about the manipulated object’s pose and stability, but also of the other elements in the scene. Consequently, a key challenge in interactive perception is to select actions carefully, e.g., guided by uncertainty.
@@
To tackle this challenge, predictive networks [8, 9] have been proposed to quantify the effect of pushing objects in table-top clutter for object retrieval and action selection between grasping and pushing. However, these methods act in perceptually easy arrangements and do not consider uncertainty-awareness. Recently, Marques et al. [10] demonstrated that evidential networks can significantly improve Manipulation-Enhanced Mapping (MEM) in confined shelf spaces by integrating observation and manipulation networks that jointly predict post-action occupancy and semantic map beliefs. However, they do not leverage belief uncertainty for continuous action planning, thereby presenting several opportunities for performance improvement. Firstly, instead of fixed view points, continuous-space next-best view planning enables fine-grained sensor placements that reveal occluded regions more effectively. Secondly, instead of using a selection strategy that randomly samples push candidates, informed sampling can be used to select pushes based on their expected impact on hidden areas, improving planning efficiency and interaction safety. Finally, driving exploration with semantic uncertainty rather than occupancy entropy alone can improve both occupancy estimates and semantic scene understanding.
Figure 1: The task of the robot is to map a shelf with dense clutter and heavy occlusions with its RGB-D sensor mounted on the arm. Highly uncertain areas are marked in white in the belief map whereas solid colors indicate low semantic uncertainty regions. Initially, the arm finds the next-best view targeting the highly uncertain regions on the right. Once frontier objects are mapped, the belief map indicates uncertain regions behind the occluding Schoko cereal box and Cheezit box. The informed push sampler finds the target region to uncover (red dot) and identifies the cereal box as the best candidate for occlusion removal. The green arrow indicates the chosen push candidate to unveil the region.
@@
To tackle these challenges, in this paper, we build upon [10] and propose informed strategies for action sampling and selection in the MEM-task. In particular, we introduce a reinforcement learning (RL)-based continuous viewpoint planner that uses occupancy and semantic uncertainty estimates from the predicted map to allow the agent to select next-best views that maximize expected information gain, enabling more efficient and targeted exploration. RL provides the flexibility to adapt to diverse, dynamic scenes while effectively balancing exploration and exploitation, capabilities that are difficult to achieve with heuristic methods, sampling-based planners, or POMDPs in high-dimensional spaces.
In addition, we present a novel uncertainty-informed push selection strategy that identifies objects with high potential to reduce occlusions in uncertain map regions. By predicting and selecting minimally invasive, high-impact manipulations, our system improves visibility while minimizing unnecessary or unsafe interactions.
Our experiments demonstrate that the proposed MEM planner outperforms prior approaches [10] in terms of map accuracy and action efficiency. Specifically, our system achieves higher semantic and occupancy mean Intersection-over-Union (mIoU) scores, reduces the number of required push actions, and lowers the total object displacement—while being significantly faster in planning both views and manipulations.
@@ II-1 Interactive Perception
Recent interactive perception frameworks that integrate physical interactions into the perception loop have demonstrated significant improvements in robustness and accuracy over active perception alone. Murali et al. showed up to a 36% increase in object‐pose estimation accuracy compared to an active‐vision baseline in dense clutter [11]. Similarly, Yao et al. demonstrated that interactively manipulating leaves to infer the shape and position of fruits significantly improves the accuracy of these estimates [6]. Furthermore, interactive perception enables the autonomous, online construction of affordance‐based scene representations without requiring predefined object models, allowing robots to adaptively learn and operate in open, dynamic environments [5]. However, these methods do not consider the problem of dense clutter and confined spaces and approaches to uncover them.
@@ II-2 Uncertainty-Aware Action Selection
Uncertainty-aware action selection methods employ probabilistic models to inform perception and manipulation, yielding more efficient and reliable task execution than heuristic-driven strategies. ActNeRF [12] uses an ensemble of NeRF models to quantify 3D reconstruction uncertainty and choose the next-best visual or reorientation action, accelerating object modeling under occlusions compared to random or greedy approaches. Furthermore, for mobile manipulation in unstructured settings, uncertainty can be used to adaptively select between sensing and manipulation actions that minimize environmental and sensory uncertainty [13]. Uncertainty-aware manipulation planning frameworks that represent object pose as probability distributions can systematically sequence pushes, placements, and grasps to eliminate pose ambiguity, achieving higher assembly success rates than deterministic planners [14].
Recently, evidential neural networks [15] have been increasingly deployed to estimate in a single forward pass, both aleatoric and epistemic uncertainties, guiding action selection in perception and manipulation. vMF-Contact [16] applies evidential learning with von Mises–Fisher models for grasping in cluttered environments, while Durasov et al. [17] leverages a Dirichlet-based evidential loss on bird’s-eye-view representations for 3D object detection to identify out-of-distribution objects and boost detection performance.
@@ II-3 Manipulation‐Enhanced Mapping
Robotic mapping in densely cluttered environments requires strategic perception, deciding where to observe, and proactive interaction, deciding how to manipulate occluding objects. Classical mechanical‐search approaches leverage semantic relationships to guide push‐and‐pick actions for target retrieval [18, 3, 19]. Interactive mapping paradigms extend this idea, using manipulation primitives to resolve occlusions and complete spatial reconstructions. Wang et al. [20] build an actionable 3D relational object graph to perform targeted interactions to explore hidden regions in mobile settings. In the domain of cluttered confined spaces, such as shelves, Dengler et al. [4] proposed a deep RL framework that alternates viewpoint planning with push‐action sampling on saturation of the marginal map entropy reduction. More recently, Marques et al. [10] adapted this idea and defined a POMDP solver whose belief resides in a metric‐semantic grid map, with evidential neural‐accelerated belief updates to reason about object shapes and manipulation consequences, achieving substantial gains in map accuracy.
Our work builds on these lines of research by leveraging the uncertainty estimates for occupancy and semantic prediction modeled using evidential networks [15] for learned continuous‐space viewpoint planning and informed occlusion‐critical push sampling, rather than random sampling or push‐scoring networks [4]. Compared to prior MEM [10], our method concentrates exploration on regions with maximum expected information gain, reduces redundant manipulations, and lowers planning overhead—enabling online semantic mapping in cluttered, dynamic scenes.
@@ III Our Approach
In this work, we address manipulation‐enhanced mapping (MEM) of cluttered shelf environments using a robotic arm with a wrist‐mounted RGB-D sensor. An illustration of our pipeline is shown in Fig. 2. Building on the evidential metric-semantic mapping of Marques et al. [10] via Calibrated Neural-Accelerated Belief Update (CNABU) networks, we fuse depth and RGB-D images into occupancy (Beta) and semantic (Dirichlet) maps, where each belief distribution inherently encodes both the predicted outcome and the associated uncertainty. An next-best view RL-policy (orange) then selects $\mathrm{SE}(3)$ camera poses to observe regions of high-uncertainty. As pure active sensing exhibits submodular returns, when its expected volumetric information gain falls below a threshold, an uncertainty‐informed push selection module (blue) identifies occlusion‐critical objects via distance and visibility transforms, samples minimally invasive pushes, and evaluates their information gain. Finally, the action selection module (grey) chooses between sensing and manipulation based on their respective expected information gains, iterating until the mapping objective is achieved. The individual parts of our system are now introduced in detail.
@@ III-A Map Belief Representation
Accurate decision‐making in MEM‐tasks relies on reasoning about both observations and confidence. Following the CNABU framework [10], we represent the map belief as a probabilistic estimate over occupancy and semantic categories, enriched with explicit uncertainty. CNABU networks are designed to perform belief propagation in metric-semantic maps by predicting both semantic and occupancy beliefs, along with well-calibrated uncertainty estimates. They incorporate prior knowledge about object geometry, occlusions, and manipulation dynamics, enabling reliable forecasting of scene changes due to both observations and physical interactions. The occupancy belief is a 3D voxel map of size $H\times W\times D$, where each voxel is parameterized by $\bm{\lambda}^{O}=(\alpha,\beta)$ and modeled by a Beta distribution whose predictive mean $\mathbb{E}[\mathrm{Beta}(\alpha,\beta)]$ is the expected occupancy probability and whose variance
$u_{o}=\mathrm{Var}[\mathrm{Beta}(\alpha,\beta)]$
quantifies epistemic uncertainty. The semantic belief is a 2D top‐down projected map of size $H\times W$ with evidential parameters $\bm{\lambda}^{S}\in\mathds{R}^{H\times W\times N_{\mathrm{classes}}}$, where $\bm{\lambda}^{S}_{i,j}=(\lambda_{1},\dots,\lambda_{N})$ defines a Dirichlet distribution whose predictive class probabilities
$\mathbb{E}[\mathrm{Dir}(\lambda_{1},\dots,\lambda_{N})]_{n}=\frac{\lambda_{n}}{S},\quad S=\sum_{n=1}^{N}\lambda_{n},$
yield a hard label $\arg\max_{n}(\,\frac{\lambda_{n}}{S})$, and whose uncertainty
$u_{s}=\frac{N}{S}$
highlights low‐confidence regions [15]. In our framework, both $u_{o}$ and $u_{s}$ guide action selection to maximize expected information gain.
@@ III-B Next-Best View Selection
To achieve effective scene coverage, we integrate the predictive capabilities of CNABU networks to estimate the map for the next time step [10] with a reinforcement learning (RL) agent that operates directly on the resulting evidential map representations. Our RL-agent receives as input both the occupancy and semantic beliefs, along with their corresponding uncertainty estimates. By explicitly leveraging these uncertainty measures, the agent is guided toward viewpoints that are expected to maximize information gain. In the following, we detail the architecture and training procedure of the RL-based viewpoint planner.
@@ III-B1 Action
We define the action space as a 6D vector $(x_{cam},y_{cam},z_{cam},target_{x},target_{y},target_{z})$, with each component normalized to $[-1,1]$. At execution time these values are rescaled to predefined physical limits in SE(3). We represent orientation implicitly via a look-at target to ensure the camera faces the region of interest (RoI), i.e., the shelf board.
@@ III-B2 Observation
(i) Enhanced world representation: As described in Sec. III-A, the agent’s world representation comprises of the 3D occupancy and 2D semantic belief maps. We project the thresholded occupancy belief into a 2D maximum-height map to reduce memory and computation while preserving occlusion reasoning and visibility assessment. Additionally, we incorporate cell-wise occupancy and semantic uncertainties, i.e., $u_{o}$ and $u_{s}$, from the Beta and Dirichlet distributions to enable the agent to learn how its actions affect scene uncertainty.
(ii) Action history awareness: To avoid redundant observations and encourage coverage diversity, the observation space includes the last $N_{\mathrm{hist}}$ executed actions along with their corresponding 2D height map observations. This temporal context helps the agent identify previously explored areas and prioritize novel or more informative viewpoints.
TABLE I: Observation space used by the RL agent.
@@ III-B3 Reward
The reward function comprises three terms:
$r=\begin{cases}-r_{\text{feasibility}},&\text{if unfeasible}\\ \alpha\cdot r_{\text{uncertainty}}+\beta\cdot r_{\text{repeat}},&\text{otherwise}\end{cases}$ (1)
@@
Here, $p_{\text{dist}}$ and $e_{\text{dist}}$ denote the Euclidean distance between positions and the cosine distance between the rotation vectors, respectively, within the action history stack $\mathcal{H}$. The terms $\gamma_{p}$ and $\gamma_{e}$ are scaling factors, while $\theta_{p}$ and $\theta_{e}$ are the corresponding distance thresholds. The repetition penalty $r_{\text{repeat}}$ is computed only for entries in $\mathcal{H}$ where $p_{\text{dist}}\leq\theta_{p}$ and $e_{\text{dist}}\leq\theta_{e}$.

### v2

@@ Abstract
Service robots operating in cluttered human environments such as homes, offices, and schools cannot rely on predefined object arrangements and must continuously update their semantic and spatial estimates while dealing with possible frequent rearrangements. Efficient and accurate mapping under such conditions demands selecting informative viewpoints and targeted manipulations to reduce occlusions and uncertainty. In this work, we present a manipulation-enhanced semantic mapping framework for occlusion-heavy shelf scenes that integrates evidential metric-semantic mapping with reinforcement-learning-based next-best view planning and targeted action selection. Our method thereby exploits uncertainty estimates from Dirichlet and Beta distributions in the map prediction networks to guide both active sensor placement and object manipulation, focusing on areas with high uncertainty and selecting actions with high expected information gain. Furthermore, we introduce an uncertainty-informed push strategy that targets occlusion-critical objects and generates minimally invasive actions to reveal hidden regions by reducing overall uncertainty in the scene. The experimental evaluation shows that our framework enables to accurately map cluttered scenes, while substantially reducing object displacement and achieving a 95% reduction in planning time compared to the state-of-the-art, thereby realizing real-world applicability.
@@ I Introduction
The successful deployment of general-purpose service robots in homes and offices relies on perceiving and manipulating diverse objects in cluttered and constrained spaces. Robots must move beyond passive perception (e.g., mapping with a static camera) and instead apply active perception [1] in combination with deliberate physical object interactions [2]. Interactive perception techniques can remove occlusions by revealing hidden regions [3, 4], which is crucial for exploring unknown environments [5], building detailed object models [6], and locating occluded objects [7]. However, the involved manipulation actions come at a cost, as uninformed interactions can increase uncertainty in the target object’s pose and stability, as well as in the neighboring scene elements. Hence, a central challenge in interactive perception is uncertainty-aware action selection.
@@
To address this challenge, predictive networks have been used to estimate the effects of pushes in tabletop clutter for object retrieval and grasp–push selection [8, 9], but they target perceptually simple setups and ignore uncertainty. Marques et al. [10] showed that evidential networks improve Manipulation-Enhanced Mapping (MEM) in confined shelves by jointly predicting post-action occupancy and semantic beliefs. However, they do not leverage belief uncertainty for continuous action planning. Therefore, we propose three improvements: (i) continuous-space next-best-view (NBV) planning for fine-grained sensor placement that better reveals occlusions; (ii) uncertainty-informed push sampling that selects candidates by expected impact on hidden regions, improving efficiency and safety, i.e., reducing collisions with other objects; and (iii) NBV exploration driven by semantic uncertainty rather than occupancy entropy alone, improving both occupancy and semantic beliefs.
Figure 1: The robot employs its arm‑mounted RGB‑D sensor to map a densely cluttered shelf. In the belief map (bottom), white denotes high uncertainty, while solid colors show confidently mapped areas. The robot first uses a next‑best‑view strategy to target uncertain regions on the right of the shelf. Once those frontier objects are mapped, uncertain zones remain behind the cereal and Cheezit boxes. Our informed push sampler then most informative uncertain point (red dot), selects the cereal box since it occludes the area behind it, and pushes it (green arrow) to reveal the occluded space.
@@
We build upon [10] and propose informed strategies for action sampling and selection in the MEM-task. In particular, we propose a Reinforcement Learning (RL) based continuous NBV planner that uses predicted occupancy and semantic uncertainty to maximize expected information gain, enabling targeted exploration. RL enables the NBV planner to adapt to diverse, dynamic scenes and balances exploration–exploitation without crafted heuristics or high-dimensional POMDPs.
Furthermore, we introduce uncertainty-informed push selection that targets occluding objects predicted to reduce uncertainty, choosing minimally invasive, high-impact manipulations to improve visibility of hidden regions. Fig. 1 illustrates the two stages of our planning framework: First, the execution of a next‑best‑view followed by an uncertainty‑informed push to expose a highly uncertain region behind two view-blocking boxes.
Our experiments demonstrate that our novel uncertainty-informed MEM planner accurately maps cluttered scenes and outperforms the prior approach [10] in terms of action efficiency, i.e., it reduces the number of required push actions, and lowers the total number of object displacement, while being significantly faster in planning time. An implementation of our method can be found on Github [^1] https://github.com/HumanoidsBonn/manipulation_enhanced_map_prediction.
@@ II-1 Interactive Perception
Recent interactive perception frameworks that integrate physical interactions into the perception loop have demonstrated significant improvements in robustness and accuracy over active perception alone. For example, Murali et al. report huge increase in object‐pose estimation accuracy compared to a baseline in dense clutter [11]. Similarly, Yao et al. demonstrated that manipulating leaves to reveal fruit markedly improves shape and pose estimates [6]. Furthermore, interactive perception enables the autonomous, online construction of affordance‐based scene representations without requiring predefined object models [5]. However, these methods do not consider the problem of dense clutter at all,or in confined spaces, and approaches to uncover them.
@@ II-2 Uncertainty-Aware Action Selection
Uncertainty-aware action selection methods employ probabilistic models to inform perception and manipulation, yielding more efficient and reliable task execution than heuristic-driven strategies. For example, ActNeRF [12] quantifies 3D reconstruction uncertainty with NeRF ensembles to choose the next-best view or reorientation, accelerating modeling under occlusions versus random/greedy baselines. Furthermore, for mobile manipulation, uncertainty can arbitrate between sensing and acting to reduce environmental and sensory uncertainty [13]. Representing object pose as a distribution also enables assembly planners to sequence pushes, placements, and grasps to eliminate ambiguity, attaining higher success than deterministic methods [14]. Additionally, evidential neural networks [15] provide single-pass estimates of aleatoric and epistemic uncertainty for action selection. vMF-Contact [16] applies evidential vMF models for grasping in clutter, while Durasov et al. [17] use Dirichlet-based losses on 2D bird’s eye view representations for 3D detection to identify out of distribution objects and improve detection.
@@ II-3 Manipulation‐Enhanced Mapping
Robotic mapping in dense clutter requires strategic perception (where to observe) and proactive interaction (how to manipulate occluding objects). Classical mechanical-search approaches exploit semantic relations to guide push-and-pick actions for target retrieval [18, 3, 19]. Interactive mapping extends this idea with manipulation primitives to resolve occlusions. Wang et al. [20] use an actionable 3D relational object graph for targeted interactions in mobile settings, whereas Dengler et al. [4] alternate next-best-view planning with push sampling once marginal map-entropy reduction saturates, for mapping of cluttered shelves. More recently, Marques et al. [10] adapted this idea and defined a POMDP solver whose belief resides in a metric‐semantic grid map, with evidential calibrated neural‐accelerated belief updates (CNABU) to reason about object shapes and manipulation consequences, leading to higher map accuracy.
Building on [10], we retain its evidential belief predictors and explicitly exploit their occupancy and semantic uncertainty to learn continuous-space NBV planning and occlusion-critical push sampling. This extends MEM from fixed viewpoints and random, or scoring-based pushes to uncertainty-guided actions that focus on high information-gain regions, reduce redundant manipulation, and lower planning cost, thus enabling online semantic mapping in cluttered, dynamic scenes.
@@ III Our Approach
In this work, we address manipulation‐enhanced mapping (MEM) of cluttered shelf environments using a robotic arm with a wrist‐mounted RGB-D sensor. An illustration of our pipeline is shown in Fig. 2. Building on evidential metric-semantic mapping of Marques et al. [10] via CNABU networks, we fuse RGB-D into occupancy (Beta) and semantic (Dirichlet) belief maps that encode predictions and their uncertainty. A next-best view RL-policy (orange) then selects $\mathrm{SE}(3)$ camera poses to observe regions of high-uncertainty. As active sensing shows diminishing returns, once the expected volumetric information gain drops below a threshold, an uncertainty-informed push selector (blue) identifies occlusion-critical objects using distance and visibility transforms, samples minimally invasive pushes, and estimates their information gain. An action selector (grey) then chooses sensing or manipulation by comparing expected gains, iterating until the mapping objective is met. The individual parts of our system are now introduced in detail.
@@ III-A Map Belief Representation
Accurate decision‐making in MEM‐tasks relies on reasoning about both observations and confidence. Following the CNABU framework [10], we represent the map belief as a probabilistic estimate over occupancy and semantic categories, enriched with explicit uncertainty. CNABU networks are designed to propagate belief in metric-semantic maps by predicting both semantic and occupancy beliefs, along with well-calibrated uncertainty estimates. They use Beta and Dirichlet distributions as conjugate priors for Bernoulli (occupancy) and multinomial (semantic segmentation) likelihoods, respectively, ensuring that the resulting posteriors remain within the same distribution family. This property allows for closed-form estimation of expected values and predictive uncertainty, avoiding the computational overhead of sampling-based methods such as Monte Carlo dropout or deep ensembles. Furthermore, they incorporate prior knowledge about object geometry, occlusions, and manipulation dynamics, enabling reliable forecasting of scene changes due to both observations and physical interactions.
The occupancy belief $O$ is a 3D voxel map of size $H\times W\times D$, where each voxel is parameterized by $\bm{\lambda}^{O}=(\alpha,\beta)$ and modeled by a Beta distribution whose predictive mean $\mathbb{E}[\mathrm{Beta}(\alpha,\beta)]$ is the expected occupancy probability and whose variance $u_{o}=\mathrm{Var}[\mathrm{Beta}(\alpha,\beta)]$ quantifies epistemic uncertainty. The semantic belief $S$ is a 2D top‐down projected map of size $H\times W$ with evidential parameters $\bm{\lambda}^{S}\in\mathds{R}^{H\times W\times N_{\mathrm{classes}}}$, where $\bm{\lambda}^{S}=(\lambda_{1},\dots,\lambda_{N})$ defines a Dirichlet distribution whose predictive class probabilities
$\mathbb{E}[\mathrm{Dir}(\lambda_{1},\dots,\lambda_{N})]_{n}=\frac{\lambda_{n}}{S},\quad S=\sum_{n=1}^{N}\lambda_{n},$ (1)
yield a hard label $\arg\max_{n}(\,\frac{\lambda_{n}}{S})$, and whose uncertainty $u_{s}=N/S$ highlights low‐confidence regions [15]. In our framework, both $u_{o}$ and $u_{s}$ guide action selection to maximize expected information gain.
@@ III-B Next-Best View Selection
To achieve efficient scene coverage, we couple CNABU’s one-step predictive mapping [10] with a reinforcement learning (RL) agent that operates directly on the evidential map representations. The agent receives occupancy and semantic beliefs together with their uncertainties and selects $\mathrm{SE}(3)$ viewpoints to maximize expected information gain, guiding it toward the most informative, occluded regions.
@@ III-B1 Action
We define the action space as a 6D vector $(x_{cam},y_{cam},z_{cam},x_{target},y_{target},z_{target})$, with each component normalized to [1,1][-1,1]. At execution time these values are rescaled to predefined physical limits in SE(3). In this action space, camcam represents the desired 3D position of the camera. Furthermore, we represent orientation implicitly via a look-at targettarget to ensure the camera faces the region of interest (RoI), i.e., the shelf board.
@@ III-B2 Observation
(i) Enhanced world representation: As described in Sec. III-A, the agent’s world representation comprises of the 3D occupancy and 2D semantic belief maps of known size. We project the thresholded occupancy belief into a 2D maximum-height map to reduce memory and computation while preserving occlusion reasoning and visibility assessment. Additionally, we incorporate cell-wise occupancy and semantic uncertainties, i.e., $u_{o}$ and $u_{s}$, from the Beta and Dirichlet distributions to enable the agent to learn how its actions affect scene uncertainty.
(ii) Action history awareness: To avoid redundant viewpoints and encourage coverage diversity, the observation space includes the last $N_{\mathrm{hist}}$ executed actions along with their corresponding 2D height map observations. This temporal context helps the agent identify previously explored areas and prioritize novel or more informative viewpoints.
@@ III-B3 Reward
The reward function comprises three terms with $w_{1}$ and $w_{2}$ as scaling factors:
$r=\begin{cases}-r_{\text{feasibility}},&\text{if unfeasible}\\ w_{1}\cdot r_{\text{uncertainty}}+w_{2}\cdot r_{\text{repeat}},&\text{otherwise}\end{cases}\vskip-10.00002pt$ (2)
@@
Here, $p_{\text{dist}}$ and $e_{\text{dist}}$ denote the Euclidean distance between positions and the cosine distance between the rotation vectors, respectively, within $\mathcal{H}$. The terms $\gamma_{p}$ and $\gamma_{e}$ are scaling factors, while $\theta_{p}$ and $\theta_{e}$ are the corresponding distance thresholds. The repetition penalty $r_{\text{repeat}}$ is computed only for entries in $\mathcal{H}$ where $p_{\text{dist}}\leq\theta_{p}$ and $e_{\text{dist}}\leq\theta_{e}$.